OpenStudy (gabylovesyou):

What is the value of x rounded to the nearest hundredth? Type the numeric answer only in the box below.

OpenStudy (anonymous):

We need more information

OpenStudy (anonymous):

do you have the problem?

OpenStudy (gabylovesyou):



OpenStudy (gabylovesyou):

@mathslover @elementwielder

mathslover (mathslover):

Use tan theta = \(\cfrac{opposite}{adjacent}\) So, \(\tan 49 = \cfrac{x}{55}\)

OpenStudy (anonymous):

do you know sin cos and tan?

OpenStudy (anonymous):

@mathslover got it above ^

OpenStudy (gabylovesyou):

x = 2695 ? o-o @mathslover (sorry i suck at geometry) lol

OpenStudy (gabylovesyou):

@elementwielder yes im looking at the equations now

mathslover (mathslover):

@Gabylovesyou can you tell me what is tan 49?

mathslover (mathslover):

Use your calculator for that.

OpenStudy (gabylovesyou):

1.15

OpenStudy (anonymous):

yes

OpenStudy (gabylovesyou):

yay xD

mathslover (mathslover):

Fine. So, 1.15 = \(\cfrac{x}{55}\) Multiply 55 both sides : \(1.15 \times 55 = \cfrac{x}{55} \times 55\) \(1.15 \times 55 = \cfrac{x}{\cancel{55}} \times \cancel{55}\)

OpenStudy (anonymous):

x=?

OpenStudy (gabylovesyou):

ohhhhhh 63.25

OpenStudy (anonymous):

mhmm

mathslover (mathslover):

Yes. Good work @Gabylovesyou

OpenStudy (gabylovesyou):

ty... it says rounded to the nearest hundred.. i dont have to do anything to it right ?

OpenStudy (anonymous):

no it is good

OpenStudy (gabylovesyou):

thanks!

OpenStudy (anonymous):

yw anytime